# Improving Accuracy and Generalization for Efficient Visual Tracking

### v1

@@ Abstract
Efficient visual trackers overfit to their training distributions and lack generalization abilities, resulting in them performing well on their respective in-distribution (ID) test sets and not as well on out-of-distribution (OOD) sequences, imposing limitations to their deployment in-the-wild under constrained resources. We introduce SiamABC, a highly efficient Siamese tracker that significantly improves tracking performance, even on OOD sequences. SiamABC takes advantage of new architectural designs in the way it bridges the dynamic variability of the target, and of new losses for training. Also, it directly addresses OOD tracking generalization by including a fast backward-free dynamic test-time adaptation method that continuously adapts the model according to the dynamic visual changes of the target. Our extensive experiments suggest that SiamABC shows remarkable performance gains in OOD sets while maintaining accurate performance on the ID benchmarks. SiamABC outperforms MixFormerV2-S by 7.6% on the OOD AVisT benchmark while being 3x faster (100 FPS) on a CPU.
@@ 3.1 Architecture
The dual-template features and the dual-search-region features used by the correlation module $CC(\cdot,\cdot)$ could each simply be the naive concatenation of the respective backbone features in each pair. However, we can potentially improve their combination by processing them with an efficient learnable layer that filters out less useful components while enhancing those important for the task at hand. This should hopefully lead to improved performance. In Figure 4(top), we have shown just that. Especially when tracking in the OOD case, naive feature concatenation severely underperforms the filtered combination.
@@ 3.2 Training Losses
where $||\cdot||_{2}$ is the $\ell_{2}$-norm, $h_{1}$ and $h_{2}$ are MLP projection heads used only during training, and $D(\cdot,\cdot)$ calculates the cosine distance. Moreover, $\mathcal{D}(\cdot,\cdot)$ is computed as in [7], where we implement the stop-gradient operation to avoid degenerated solutions. We refer to the pair $\mathcal{L}_{TR}$ and $\mathcal{L}_{Reg}$ as the transitive relation loss (TRL) since it is meant to bridge the similarities between template and search region images with the aid of the dynamic components so that the relevant relational differences can be highlighted by the downstream blocks for task purposes. The TRL loss improves tracking performance, especially when both components are used. See Figure 4(middle). Notably, the AUC on AVisT [45] improves from 43.7% to 45.8%.
@@
$\mathcal{L}=\lambda_{IoU}\mathcal{L}_{IoU}+\lambda_{FL}\mathcal{L}_{FL}+\lambda_{TR}\mathcal{L}_{TR}+\lambda_{Reg}\mathcal{L}_{Reg},$ (3)
where we set $\lambda_{IoU}$, $\lambda_{FL}$, $\lambda_{TR}$, and $\lambda_{Reg}$ to 1.5, 1.5, 0.5, and 0.5, respectively.
@@
where, ${\rho}_{t}$ is the score at time $t$, and ${\lambda}_{D}$ is a momentum parameter, which we set to 0.25. We also start a counter, $C$, that when it reaches, let us say $N=60$ frames ($\approx$ 2 seconds), we compare the current classification score with the running average, and if ${\rho}_{t}>\overline{\rho}_{t-1}$, then we update the dynamic image components and reset the counter, otherwise we repeat the test at the next iteration. This strategy is parameterless, uses limited computational resources, and is effective.
@@ 3.3 Dynamic Test-Time Adaptation
Prior works have explored BN adaptation for classification purposes, with some using a backward pass for adaptation [44, 52], while others introduced backward-free adaptation [46, 49, 41, 33]. However, none are truly applicable for tracking. As we show in Table 5, applying a new Instance-Norm (IN) layer [46] is expensive, and slows down the speed twofold without noticeable improvement. Additionally, the batch size for tracking remains 1, which is too small to make any significant improvement with [49] that applies weighted momentum based on the target batch-size. Given the scale of our architecture and limited number of parameters, by replacing batch statistics with instance statistics, AdaBN [33] does not improve the performance either. DUA [41] uses source statistics as a prior for the incoming task but does not stay anchored to the source statistics, resulting in target statistics drifting away from the original distribution, causing performance drop. Therefore, we propose that BN statistics should be updated with weighted instance statistics while remaining anchored to the source statistics. This results in the following strategy
@@
HiT-Small [24]  [^1] HiT [24] could not release their code and the performance was not ascertained; therefore, we report Hit-Small from their publication as the most comparable tracker. According to those numbers, we outperform Hit-Small in various benchmarks as shown in Table 3. - - - 0.618 - 0.633 - 0.777 0.819 0.731 0.626 0.712 0.605 0.615 - - -
@@ 4.2 Comparison with other Trackers
AVisT[45] consists of 120 extremely challenging real-world sequences in adverse visibility (see Figure 2 in the supplementary material). In addition to simple occlusion and fast motion, it involves objects under heavy rain, heavy snow, dense fog, sandstorms, hurricanes, etc. In Table 3, we note that this out-of-distribution benchmark highlights a significant performance degradation of SOTA trackers compared to the widely used in-distribution test sets. On the other hand, our trackers, S-Tiny and S-Small, are outperforming HCAT by 5.4% and 6.1% respectively and MixFormerV2-S[12] by 7.6% and 8.3% respectively, showing our approach’s improved ability to track in sequences with adverse conditions. Additionally, S-Tiny outperforms SMAT[18] by 2.5% while being almost 3x faster on a CPU.
@@ 4.3 Comparision with Adaptation Approaches
In our S-Tiny model we incorporate TENT[52] and ETA[44] as our TTA baselines with backward-passes, where we use our classification output as self-entropy. Additionally, we also evaluated Momentum[49], DUA[41], IN[46], and AdaBN[33] as our backward-free TTA baselines. We set the batch size to 1. The comparison is shown in Table 5 with two of the most challenging benchmarks, AVisT as it involves multiple adverse scenarios with natural corruptions and ITB as the next most challenging benchmark. TENT and ETA have almost 3x the latency because of the backward basses, and do improve with AVisT. However, we do not observe the same with ITB. Momentum shows negligible improvement, whereas IN, AdaBN, and DUA show performance degradation in both scenarios. When DTTA, our efficient adaptation strategy, is turned on, we notice significant improvement with both benchmarks while having minimal latency.

### v2

@@ Abstract
Efficient visual trackers overfit to their training distributions and lack generalization abilities, resulting in them performing well on their respective in-distribution (ID) test sets and not as well on out-of-distribution (OOD) sequences, imposing limitations to their deployment in-the-wild under constrained resources. We introduce SiamABC, a highly efficient Siamese tracker that significantly improves tracking performance, even on OOD sequences. SiamABC takes advantage of new architectural designs in the way it bridges the dynamic variability of the target, and of new losses for training. Also, it directly addresses OOD tracking generalization by including a fast backward-free dynamic test-time adaptation method that continuously adapts the model according to the dynamic visual changes of the target. Our extensive experiments suggest that SiamABC shows remarkable performance gains in OOD sets while maintaining accurate performance on the ID benchmarks. SiamABC outperforms MixFormerV2-S by 7.6% on the OOD AVisT benchmark while being 3x faster (100 FPS) on a CPU. Our code and models are available at https://wvuvl.github.io/SiamABC/.
@@ 3.1 Architecture
The dual-template features and the dual-search-region features used by the correlation module $CC(\cdot,\cdot)$ could each be the naive concatenation of the respective backbone features in each pair. However, we can potentially improve their combination by processing them with an efficient learnable layer that filters out less useful components while enhancing those important for the task at hand. This should hopefully lead to improved performance. In Figure 4(top), we have shown just that. Especially when tracking in the OOD case, naive feature concatenation severely underperforms the filtered combination.
@@ 3.2 Training Losses
where $||\cdot||_{2}$ is the $\ell_{2}$-norm, $h_{1}$ and $h_{2}$ are MLP projection heads used only during training, and $D(\cdot,\cdot)$ calculates the cosine distance. Moreover, $\mathcal{D}(\cdot,\cdot)$ is computed as in [7], where we implement the stop-gradient operation to avoid degenerated solutions. We refer to the pair $\mathcal{L}_{TR}$ and $\mathcal{L}_{Reg}$ as the transitive relation loss (TRL) since it is meant to bridge the similarities between template and search region images with the aid of the dynamic components so that the relevant relational differences can be highlighted by the downstream blocks for task purposes. The TRL loss improves tracking performance, especially when both components are used. See Figure 4(middle). Notably, in the OOD case, the AVisT [45] AUC improves from 43.7% to 45.8%.
@@
$\small\centering\mathcal{L}=\mathcal{L}_{IoU}+\lambda_{FL}\mathcal{L}_{FL}+\lambda_{TR}\mathcal{L}_{TR}+\lambda_{Reg}\mathcal{L}_{Reg},\@add@centering$ (3)
where we set $\lambda_{FL}$, $\lambda_{TR}$, and $\lambda_{Reg}$ to 1, 1/3, and 1/3, respectively.
@@
where, ${\rho}_{t}$ is the score at time $t$, and ${\lambda}_{D}$ is a momentum parameter, which we set to 0.25. We also start a counter, $C$, that when it reaches, let us say $N=60$ frames ($\approx$ 2 seconds), we compare the current classification score with the running average, and if ${\rho}_{t}>\overline{\rho}_{t-1}$, then we update the dynamic image components and reset the counter, otherwise we repeat the test at the next iteration. This strategy is effective, parameterless, and uses limited computational resources.
@@ 3.3 Dynamic Test-Time Adaptation
Prior works have explored BN adaptation for classification purposes, with some using a backward pass for adaptation [44, 52], while others introduced backward-free adaptation [46, 49, 41, 33]. However, none are directly applicable for tracking. As we show in Table 5, applying a new Instance-Norm (IN) layer [46] is expensive, and slows down the speed twofold without noticeable improvement. Additionally, the batch size for tracking remains 1, which is too small to make any significant improvement with [49] that applies weighted momentum based on the target batch-size. Given the scale of our architecture and limited number of parameters, by replacing batch statistics with instance statistics, AdaBN [33] does not improve the performance either. DUA [41] uses source statistics as a prior for the incoming task but does not stay anchored to the source statistics, resulting in target statistics drifting away from the original distribution, causing performance drop. Therefore, we propose that BN statistics should be updated with weighted instance statistics while remaining anchored to the source statistics. This results in the following strategy
@@
HiT-Small [24] - - - 0.618 - 0.633 - 0.777 0.819 0.731 0.626 0.712 0.605 0.615 - - -
@@ 4.1 Comparison with other Trackers
AVisT[45] consists of 120 extremely challenging real-world sequences in adverse visibility. In addition to simple occlusion and fast motion, it involves objects under heavy rain, heavy snow, dense fog, sandstorms, hurricanes, etc. In Table 3, we note that this out-of-distribution benchmark highlights a significant performance degradation of SOTA trackers compared to the widely used in-distribution test sets. On the other hand, our trackers, S-Tiny and S-Small, are outperforming HCAT by 5.4% and 6.1% respectively and MixFormerV2-S[12] by 7.6% and 8.3% respectively, showing our approach’s improved ability to track in sequences with adverse conditions. Additionally, S-Tiny outperforms SMAT[18] by 2.5% while being almost 3x faster on a CPU.
@@ 4.2 Comparision with Adaptation Approaches
In our S-Tiny model we incorporate TENT[52] and ETA[44] as our TTA baselines with backward-passes, where we use our classification output as self-entropy. Additionally, we also evaluated Momentum[49], DUA[41], IN[46], and AdaBN[33] as our backward-free TTA baselines. We set the batch size to 1. The comparison is shown in Table 5 with two of the most challenging benchmarks, AVisT as it involves multiple adverse scenarios with natural corruptions and ITB as the next most challenging benchmark. TENT and ETA have almost 3x the latency because of the backward passes, and do improve with AVisT. However, we do not observe the same with ITB. Momentum shows negligible improvement, whereas IN, AdaBN, and DUA show performance degradation in both scenarios. When DTTA, our efficient adaptation strategy, is turned on, we notice significant improvement with both benchmarks while having minimal latency.